# A uniformity principle for spatial matching

### v1

@@ 5 Conclusion
Building on this insight, we introduced a dual service-range model with inflexible and flexible supply nodes and developed a Markov-chain embedding that yields tractable characterizations and bounds for the expected matching rate, including closed-form expressions. Taken together, these results provide structural guidance for spatial platforms – such as autonomous fleets, drone delivery, emergency services, and feature-based matching markets – by showing that, when designing coverage ex ante, it is typically better to spread flexibility broadly than to concentrate it on a small set of “super” units.
Open problems.
Our analysis suggests several directions for future work. First, simulations in Section 4 indicate that the uniformity principle is tightly linked to allocating service range and can fail when the decision variable is a service radius. While radius-based controls are relevant in some applications (e.g., maximum pickup distance or response-time thresholds), range-based controls are natural in high-dimensional feature spaces and in incentive designs that expand acceptance sets without changing geography. Characterizing when a uniformity (or anti-uniformity) principle holds for radius allocations, and identifying the right budget normalization when radius and covered range diverge, remains open. Second, we study an ex ante benchmark in which ranges are chosen before supply and demand locations are realized. In many platforms, the spatial distribution of supply is relatively stable (e.g., depots or driver home bases), suggesting a two-stage design where ranges are chosen after observing supply locations but before demand arrives; quantifying the value of such partial adaptivity is an important direction. Finally, our results treat flexibility on the supply side only. Extending the analysis to two-sided flexibility under coupled budgets in geometric environments, in the spirit of [FMZ24] but with spatial locality and congestion, remains open, including whether flexibility should be concentrated on one side or shared across both, and whether any form of uniformity principle survives.
@@ B.1.1 Proof of Lemma 1
$i\triangleq\max\{k:z^{(r-1)}_{k}>y_{k}\},\qquad j\triangleq\min\{k>i:z^{(r-1)}_{k}<y_{k}\},$
@@ B.2.1 Proof of Lemma 9
where (a) uses (55); (b) uses that $s$ is unifomly distributed, so the probability measure is precisely the Lebesgue measure $\mathrm{Vol}(\cdot)$; (c) uses the definition of $B_{\ell}(\cdot)$. This concludes the proof.
@@ B.2.2 Proof of Lemma 10
    $\|y-d_{j}\|\leq\|y-d_{i}\|\leq(\ell/n)^{1/k},$
    i.e. inside $V_{d_{j}}(\mathcal{D}^{\overline{*}}_{+})$, being in any $B(d_{i},(\ell/n)^{1/k})$ implies being in $B(d_{j},(\ell/n)^{1/k}$.
@@ B.2.4 Proof of Lemma 12
Recall that trimming cells have side length $L_{T}=\frac{2k\gamma^{1/k}}{\varepsilon n^{1/k}}$, whereas pattern cells have side length $L_{P}=\frac{1}{n^{1/k}}w^{\prime}_{k,\gamma}$. In order for half the pattern cells to be fully contained in trimming cells, it suffices that $L_{T}\geq 2k\,L_{P}$, or equivalently that
@@
where the inequality uses P1 to lower bound $\mathrm{Vol}(B(x_{A},R))$ and $\mathrm{Vol}(B(x_{A},R^{\prime\prime})\setminus B(x_{A},R^{\prime}))$, and also uses the fact that $\kappa_{k}$ denotes the volume of a unit sphere in ${\mathbb{R}}^{k}$. Therefore, combining with (64), we get that if $\varepsilon\leq w_{k,\gamma}$, we have:
@@ C.1.1 Proof of Lemma 14
Optimality of the greedy algorithm for interval graphs is a classical result [Glo67, LJP81]. For completeness we present a simple proof specialized to our random geometric graph. For any matching $\mathcal{N}$ and non-negative integer $k$, let $\mathcal{N}_{k}\triangleq\{(d_{(i)},s_{i^{*}})\in\mathcal{N}:i\leq k\}$ be the restriction of $\mathcal{N}$ to the first $k$ demand nodes. We show via induction that for all $k$, there exists a maximum matching $\mathcal{M}^{*}$ such that $\mathcal{M}_{k}=\mathcal{M}^{*}_{k}$. The base case is true, since $\mathcal{M}_{0}=\mathcal{M}^{*}_{0}=\emptyset$ for all maximum matchings $\mathcal{M}^{*}$. Thus, assume $k\geq 1$ and $\mathcal{M}_{k-1}=\mathcal{M}^{*}_{k-1}$ for a maximum matching $\mathcal{M}^{*}$.
Case $1$. $d_{(k)}$ is unmatched by $\mathcal{M}$. This happens only if $d_{(k)}$ has no neighbors in $\mathsf{S}^{\prime}_{k}$, the set of unmatched supply nodes in iteration $k$ of Algorithm 1. In this case, $\mathcal{M}^{*}$ cannot match $d_{(k)}$, so $\mathcal{M}_{k}=\mathcal{M}^{*}_{k}$.
Case $2$. $d_{(k)}$ is matched to $s_{k^{*}}$ by $\mathcal{M}$. There are two subcases to consider.
  • $d_{(k)}$ is unmatched by $\mathcal{M}^{*}$. It must be that for some $\ell>k$, $\mathcal{M}^{*}$ matches $d_{(\ell)}$ to $s_{k^{*}}$, otherwise $\mathcal{M}^{*}+(d_{(k)},s_{k^{*}})$ would be a larger matching. Then, the matching
    $\widetilde{\mathcal{M}}\triangleq\mathcal{M}^{*}-(d_{(\ell)},s_{k^{*}})+(d_{(k)},s_{k^{*}})$
    is maximum and satisfies $\mathcal{M}_{k}=\widetilde{\mathcal{M}}_{k}$.
  • $d_{(k)}$ is matched by $\mathcal{M}^{*}$. If it is matched to $s_{k^{*}}$, then $\mathcal{M}_{k}=\mathcal{M}^{*}_{k}$ is trivially satisfied. Therefore, assume that $d_{(k)}$ is matched by $\mathcal{M}^{*}$ to some $s_{(\ell)}\neq s_{k^{*}}$. If $s_{k^{*}}$ is unmatched by $\mathcal{M}^{*}$, let $\widetilde{\mathcal{M}}$ be the matching
    $\widetilde{\mathcal{M}}\triangleq\mathcal{M}^{*}-(d_{(k)},s_{(\ell)})+(d_{(k)},s_{k^{*}}).$
    Otherwise, if $s_{(\ell)}$ is matched to some $d_{(j)}$ with $j>k$, let $\widetilde{\mathcal{M}}$ be the matching
    $\widetilde{\mathcal{M}}\triangleq\mathcal{M}^{*}-(d_{(k)},s_{(\ell)})-(d_{(j),}s_{k^{*}})+(d_{(k)},s_{k^{*}})+(d_{(j)},s_{(\ell)}).$
    To see why $(d_{(j)},s_{(\ell)})$ is a valid match, note that $d_{(k)}$ connects to both $s_{k^{*}}$ and $s_{(\ell)}$. Since $d_{(j)}$ comes after $d_{(k)}$ and connects to $s_{k^{*}}$, it must also connect to $s_{(\ell)}$ because $s_{k^{*}}+r_{k^{*}}\leq s_{(\ell)}+r_{(\ell)}$ per the matching rule. Thus, $\widetilde{\mathcal{M}}$ is a maximum matching satisfying $\mathcal{M}_{k}=\widetilde{\mathcal{M}}_{k}$.
@@ C.1.2 Proof of Lemma 15
Assume that $\mathcal{M}(\widehat{G}_{n})$ and $\widehat{\mathcal{M}}(\widehat{G}_{n})$ disagree, and let $k\geq 1$ be the smallest index at which $d_{(k)}$ is resolved differently between the two procedures. By minimality of $k$, the two runs make identical decisions for $d_{(1)},\ldots,d_{(k-1)}$; hence the sets of unmatched flexible and non‑flexible supply nodes lying to the left of $d_{(k)}$ coincide in both runs.
At the moment, the generative procedure is ready to decide on $d_{(k)}$, each supply node type $T\in\{\mathrm{F},\mathrm{NF}\}$ has been advanced until it is either in range or ahead of the demand node. Formally, letting $v^{T}$ be that type’s active supply node and $s_{T}/n$ its radius, the type status is one of:
$\text{(in range)}\quad d_{(k)}\in\bigl[v^{T}{-}s_{T}/n,\;v^{T}{+}s_{T}/n\bigr],\qquad\mbox{or}\qquad\text{(ahead)}\quad v^{T}{-}s_{T}/n>d_{(k)}.$
Because we are revealing the same PPP points left‑to‑right, the active supply node of each type in the generative run is exactly the leftmost unmatched supply node of that type in the greedy run, and its status (in‑range vs. ahead) with respect to $d_{(k)}$ is identical in both runs. There are three possibilities:
@@
    Both supply types ahead. No supply node can reach $d_{(k)}$, so both procedures leave $d_{(k)}$ unmatched.
@@
    Exactly one supply type $T$ in range. The other type is ahead and cannot serve $d_{(k)}$. Within type $T$, all supply nodes have the same range, and the leftmost unmatched feasible supply node (the active $T$ just identified) has the earliest deadline $v^{T}+s_{T}/n$ among feasible $T$-supply nodes. Hence both procedures must match $d_{(k)}$ to that same supply node.
@@
    Both supply types in range. Each procedure compares the same two deadlines $D^{\mathrm{F}}=v^{\mathrm{F}}+(r+b)/{n}$ and $D^{\mathrm{NF}}=v^{\mathrm{NF}}+r/{n},$ and, by the greedy rule, matches $d_{(k)}$ to the earlier deadline.
In all cases, the decisions coincide, which contradicts the definition of $k$. Therefore, no disagreement occurs, concluding the proof.

### v2

@@ 5 Conclusion
Building on this insight, we introduced a dual service range model with inflexible and flexible supply nodes and developed a Markov-chain embedding that yields tractable characterizations and bounds for the expected matching rate, including closed-form expressions. Taken together, these results provide structural guidance for spatial platforms such as autonomous fleets, drone delivery, emergency services, and feature-based matching markets by showing that, when designing coverage ex-ante, it is typically better to spread flexibility broadly than to concentrate it on a small set of “super” units.
Our analysis suggests several directions for future work. First, our simulations indicate that the uniformity principle depends critically on the decision variable: it holds for service range parametrized by volume but can fail for service radius parametrized by radius. Note that while these formulations are equivalent for $k=1$, they diverge for $k\geq 2$ due to the non-linear scaling of volume with radius. Characterizing optimal allocations under radius constraints, where cost and coverage volume scale non-linearly, remains an open challenge.
Second, we study an ex-ante benchmark in which ranges are chosen before supply and demand locations are realized. In many platforms, the spatial distribution of supply is relatively stable (e.g., depots or driver home bases), suggesting a two-stage design where ranges are chosen after observing supply locations but before demand arrives; quantifying the value of such partial adaptivity is an important direction.
Finally, our results treat flexibility on the supply side only. Extending the analysis to two-sided flexibility under coupled budgets in geometric environments, in the spirit of [FMZ24] but with spatial locality and congestion, remains open, including whether flexibility should be concentrated on one side or shared across both, and whether any form of uniformity principle survives.
@@ B.1.1 Proof of Lemma 1
$i\triangleq\max\{\ell:z^{(r-1)}_{\ell}>y_{\ell}\},\qquad j\triangleq\min\{\ell>i:z^{(r-1)}_{\ell}<y_{\ell}\},$
@@ B.2.1 Proof of Lemma 9
where (a) uses (56); (b) uses that $s$ is uniformly distributed, so the probability measure is precisely the Lebesgue measure $\mathrm{Vol}(\cdot)$; (c) uses the definition of $B_{\ell}(\cdot)$. This concludes the proof.
@@ B.2.2 Proof of Lemma 10
    $\|y-d_{j}\|_{2}\leq\|y-d_{i}\|_{2}\leq(\ell/n)^{1/k},$
    i.e. inside $V_{d_{j}}(\mathcal{D}^{\overline{*}}_{+})$, being in any $B(d_{i},(\ell/n)^{1/k})$ implies being in $B\left(d_{j},(\ell/n)^{1/k}\right)$.
@@ B.2.4 Proof of Lemma 12
Recall that trimming cells have side length $L_{T}={2k\gamma^{1/k}}/({\varepsilon n^{1/k}})$, whereas pattern cells have side length $L_{P}=w^{\prime}_{k,\gamma}/{n^{1/k}}$. In order for half the pattern cells to be fully contained in trimming cells, it suffices that $L_{T}\geq 2k\,L_{P}$, or equivalently that
@@
where the inequality uses P1 to lower bound $\mathrm{Vol}(B(x_{A},R))$ and $\mathrm{Vol}(B(x_{A},R^{\prime\prime})\setminus B(x_{A},R^{\prime}))$, and also uses the fact that $\kappa_{k}$ denotes the volume of a unit ball in ${\mathbb{R}}^{k}$. Therefore, combining with (65), we get that if $\varepsilon\leq w_{k,\gamma}$, we have:
@@ C.1.1 Proof of Lemma 14
Optimality of the greedy algorithm for interval graphs is a classical result [Glo67, LJP81]. For completeness we present a simple proof specialized to our random geometric graph. For any matching $\mathcal{N}$ and non-negative integer $t$, let $\mathcal{N}_{t}\triangleq\{(d_{(i)},s_{i^{*}})\in\mathcal{N}:i\leq t\}$ be the restriction of $\mathcal{N}$ to the first $t$ demand nodes. We show via induction that for all $t$, there exists a maximum matching $\mathcal{M}^{*}$ such that $\mathcal{M}_{t}=\mathcal{M}^{*}_{t}$. The base case is true, since $\mathcal{M}_{0}=\mathcal{M}^{*}_{0}=\emptyset$ for all maximum matchings $\mathcal{M}^{*}$. Thus, assume $t\geq 1$ and $\mathcal{M}_{t-1}=\mathcal{M}^{*}_{t-1}$ for a maximum matching $\mathcal{M}^{*}$.
Case $1$. $d_{(t)}$ is unmatched by $\mathcal{M}$. This happens only if $d_{(t)}$ has no neighbors in $\mathsf{S}^{\prime}_{t}$, the set of unmatched supply nodes in iteration $t$ of Algorithm 1. In this case, $\mathcal{M}^{*}$ cannot match $d_{(t)}$, so $\mathcal{M}_{t}=\mathcal{M}^{*}_{t}$.
Case $2$. $d_{(t)}$ is matched to $s_{t^{*}}$ by $\mathcal{M}$. There are two subcases to consider.
  • $d_{(t)}$ is unmatched by $\mathcal{M}^{*}$. It must be that for some $\ell>t$, $\mathcal{M}^{*}$ matches $d_{(\ell)}$ to $s_{t^{*}}$, otherwise $\mathcal{M}^{*}+(d_{(t)},s_{t^{*}})$ would be a larger matching. Then, the matching
    $\widetilde{\mathcal{M}}\triangleq\mathcal{M}^{*}-(d_{(\ell)},s_{t^{*}})+(d_{(t)},s_{t^{*}})$
    is maximum and satisfies $\mathcal{M}_{t}=\widetilde{\mathcal{M}}_{t}$.
  • $d_{(t)}$ is matched by $\mathcal{M}^{*}$. If it is matched to $s_{t^{*}}$, then $\mathcal{M}_{t}=\mathcal{M}^{*}_{t}$ is trivially satisfied. Therefore, assume that $d_{(t)}$ is matched by $\mathcal{M}^{*}$ to some $s_{(\ell)}\neq s_{t^{*}}$. If $s_{t^{*}}$ is unmatched by $\mathcal{M}^{*}$, let $\widetilde{\mathcal{M}}$ be the matching
    $\widetilde{\mathcal{M}}\triangleq\mathcal{M}^{*}-(d_{(t)},s_{(\ell)})+(d_{(t)},s_{t^{*}}).$
    Otherwise, if $s_{(\ell)}$ is matched to some $d_{(j)}$ with $j>t$, let $\widetilde{\mathcal{M}}$ be the matching
    $\widetilde{\mathcal{M}}\triangleq\mathcal{M}^{*}-(d_{(t)},s_{(\ell)})-(d_{(j),}s_{t^{*}})+(d_{(t)},s_{t^{*}})+(d_{(j)},s_{(\ell)}).$
    To see why $(d_{(j)},s_{(\ell)})$ is a valid match, note that $d_{(t)}$ connects to both $s_{t^{*}}$ and $s_{(\ell)}$. Since $d_{(j)}$ comes after $d_{(t)}$ and connects to $s_{t^{*}}$, it must also connect to $s_{(\ell)}$ because $s_{t^{*}}+r_{t^{*}}\leq s_{(\ell)}+r_{(\ell)}$ per the matching rule. Thus, $\widetilde{\mathcal{M}}$ is a maximum matching satisfying $\mathcal{M}_{t}=\widetilde{\mathcal{M}}_{t}$.
@@ C.1.2 Proof of Lemma 15
Assume that $\mathcal{M}(\widehat{G}_{n})$ and $\widehat{\mathcal{M}}(\widehat{G}_{n})$ disagree, and let $t\geq 1$ be the smallest index at which $d_{(t)}$ is resolved differently between the two procedures. By minimality of $t$, the two runs make identical decisions for $d_{(1)},\ldots,d_{(t-1)}$; hence the sets of unmatched flexible and non‑flexible supply nodes lying to the left of $d_{(t)}$ coincide in both runs.
At the moment, the generative procedure is ready to decide on $d_{(t)}$, each supply node type $T\in\{\mathrm{F},\mathrm{NF}\}$ has been advanced until it is either in range or ahead of the demand node. Formally, letting $v^{T}$ be that type’s active supply node and $s_{T}/n$ its radius, the type status is one of:
$\text{(in range)}\quad d_{(t)}\in\bigl[v^{T}{-}s_{T}/n,\;v^{T}{+}s_{T}/n\bigr],\qquad\mbox{or}\qquad\text{(ahead)}\quad v^{T}{-}s_{T}/n>d_{(t)}.$
Because we are revealing the same PPP points left‑to‑right, the active supply node of each type in the generative run is exactly the leftmost unmatched supply node of that type in the greedy run, and its status (in‑range vs. ahead) with respect to $d_{(t)}$ is identical in both runs. There are three possibilities:
@@
    Both supply types ahead. No supply node can reach $d_{(t)}$, so both procedures leave $d_{(t)}$ unmatched.
@@
    Exactly one supply type $T$ in range. The other type is ahead and cannot serve $d_{(t)}$. Within type $T$, all supply nodes have the same range, and the leftmost unmatched feasible supply node (the active $T$ just identified) has the earliest deadline $v^{T}+s_{T}/n$ among feasible $T$-supply nodes. Hence both procedures must match $d_{(t)}$ to that same supply node.
@@
    Both supply types in range. Each procedure compares the same two deadlines $D^{\mathrm{F}}=v^{\mathrm{F}}+(r+b)/{n}$ and $D^{\mathrm{NF}}=v^{\mathrm{NF}}+r/{n},$ and, by the greedy rule, matches $d_{(t)}$ to the earlier deadline.
In all cases, the decisions coincide, which contradicts the definition of $t$. Therefore, no disagreement occurs, concluding the proof.